# Local Thermal Operations and Classical Communication

### v1

@@ 4.4 Local action
Figure 6: The allowed region $\mathbf{r}=T(\mathbf{p},\mathbf{q})$ in local dimention $d=3$. The two input states are denoted by black dots. All images of $T(\mathbf{p},\mathbf{q})$ are convex combinations of extremal images, thus they lie inside the red border. The green region is the upper bound on states accessible via bi-thermal operation according to Theorem 5. (a) In infinite temperature $\beta=0$ images via permutation tensors and other extremal tensors are represented by red and blue dots, respectively. The gray dots correspond to the images of $p$ and $q$ under permutation pre-processing which does not affect the allowed region. (b) In zero temperature, $\beta\rightarrow\infty$ images via extremal tensors are denoted in red.
4.5 Action on separable states
Finally, we return to the scheme of one round of $\mathcal{S}$LTOCC between two parties to demonstrate its promising properties. Hereafter we restrict ourselves to a product input state.
@@ 4.5 Action on separable states
First and foremost we stress that $\mathcal{S}$LTOCC operation $M_{ij,kl}=T_{ijk}^{(1)}T_{jkl}^{(2)}$ is not a thermal operation. It can induce correlations even on Gibbs ensemble $\gamma=\gamma^{(1)}\otimes\gamma^{(2)}$, which can be most clearly seen in the case of extremal operations in large or small temperatures.
Lemma 2.
@@
Consider the limit of infinite $\beta\to 0$ or zero $\beta\to\infty$ temperature, and $\mathcal{S}LTOCC$ operation consisting of two copies of the same extremal bi-thermal channel $M_{ij,kl}=T_{ijk}T_{jkl}$, which entries are either 00 or $1$. Then for any product input state $\mathbf{p}\otimes\mathbf{q}$ the output state $\mathbf{\pi}=M(\mathbf{p}\otimes\mathbf{q})$ is maximally correlated.
@@
$\pi_{ij}=\sum_{kl}M_{ij,lk}p_{k}q_{l}=\sum_{kl}T_{ikl}T_{jkl}p_{k}q_{l}=\sum_{kl}\delta_{i,j}A_{ijk}p_{k}q_{l}=\delta_{ij}r_{i}\leavevmode\nobreak\ .$ (37)
Thus the output distribution is diagonal. Thus by measuring one output, one knows about the result of the second output as well. Formally, while calculating conditional entropy between outputs one obtains
$H(A|B)=H(B|A)=-\sum_{ij}\pi_{ij}\ln\left(\frac{\pi_{ij}}{r_{i}}\right)=-\sum_{ij}\delta_{i,j}r_{i}\ln\left(\frac{r_{i}}{r_{i}}\right)=-\sum_{i}r_{i}\ln(1)=0.$ (38)
The following result can be modified to encompass also some tensors in non-zero temperatures, for example, the extremal tensor at low temperatures given in (31). Notice that its last layers are identities $T_{i,j,0}=T_{i,0,k}=\mathbb{I}$. Therefore one round $\mathcal{S}$LTOCC constructed form it $M_{ij,kl}=T_{ikl}T_{jkl}$ acting on the states $\mathbf{p},\mathbf{q}$ one of which is maximally excited $\mathbf{q}=(0,\cdots,1)$ gives
$M(p,q)=\sum_{kl}T_{ikl}T_{jkl}p_{k}q_{l}=\sum_{k}T_{ik0}T_{jk0}p_{k}=\sum_{k}\delta_{i,k}\delta_{j,k}p_{k}=\delta_{i,j}p_{i}\leavevmode\nobreak\ ,$ (39)
which is once again a maximally correlated state. Moreover, its mutual information is equal to the entropy $S(p)$, so it is restricted only by $S(p)\leq\ln(n)$.
@@ 5 Summary and Outlook
In this work, we introduce a novel framework where parties operate under the distant laboratories paradigm, subject to thermodynamic constraints, coining the notion of local thermal operations and classical communication (LTOCC). We begin by presenting a hierarchy of classes of allowed operations, categorized by the level of control and the complexity of communication between the parties. A particular focus is placed on a specific subclass called symmetric LTOCC or $\mathcal{S}$LTOCC, restricted to two inputs. We explore its properties and demonstrate the complexity of this set of operations. We propose a method for generating extreme points at high temperatures by perturbing known extreme points in infinite temperature, and a concrete example of an extreme operation is provided, which exists at sufficiently low temperatures.
Next, we investigate methods to approximate accessible transformations using well-studied thermomajorization relations between the input and output states. These relations provide insight into how classical thermal constraints govern possible transformations. We also examine the impact of these operations on the joint state between the communicating parties, highlighting the high degree of correlation that emerges during the execution of the operation. This underscores the intricate interplay between thermal equilibrium and communication.
@@
The results presented here lay the foundation for a new framework that combines two established concepts: the distant laboratories setting and thermodynamic limitations on operations. Given the novelty of this framework, several key questions remain open for future research. The first is a deeper exploration of non-symmetric and multi-round scenarios, which may allow for a broader range of accessible state transitions. Another important question is the introduction of coherences into the framework. This is a challenging problem, as coherent thermal operations are difficult to describe in closed form even for a single qubit [61, 40]. The distinction between thermal operations and Gibbs-preserving operations which emerge on the quantum level further complicates this analysis.
@@
Moreover, in the face of recent developments at the frontier of quantum thermodynamics and nonlocality people have begun to investigate interplays between characteristic features of both subfields, such as possibility of generating entanglement using out-of-equilibrium states [62], detection of nonlocal properties of states using their thermodynamic properties [63] or ever-growing body of literature concerned with steady-state entanglement from autonomous machines operating in thermal framework [64, 65]. In all of the above cases, we believe that considering the problem within the LTOCC framework would further understanding of both thermodynamical and informational costs incurred throughout the implementation of the protocol.
@@
All of these questions, while significant, fall outside the scope of this initial presentation of the LTOCC framework. They should be viewed as a roadmap for future studies, aimed at expanding and refining this new theoretical approach.
@@ Acknowledgments
RB acknowledges support by the National Science Centre, Poland, under the contract number 2021/03/Y/ST2/00193 within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733.
@@ Appendix A Cooling maps
In this Appendix, we discuss the theoretical framework designed to grasp thermal operations in low temperatures: cooling maps [42]. Although the original construction was concerned with the general scenario of mixed quantum states, we let ourselves limit the discussion to diagonal states, similarly as in the rest of the paper.
@@
Therefore the energy-preserving interaction with such a bath (3) can only deexcite the system of interest. Indeed, if we consider a diagonal input state with populations $\mathbf{p}=\operatorname{diag}_{H}(\rho)$, the action of the cooling map does not create coherences and modify the populations according to [42]
@@ Appendix B Alternative derivation of cooling maps
In this Appendix we present an alternative derivation of cooling maps for diagonal states as a thermal operation in the limit of zero temperature $\beta\to\infty$.
@@
In the zero temperature limit $\beta\to\infty$, assuming nondegenerated Hamiltonian spectrum, the stochastic matrix $\Lambda$ corresponds to thermal operation if and only if $\Lambda_{i,j}=0$ for $i>j$.
The extremal operations correspond to matrices with 00 and $1$ entries, restricted only by the condition $\Lambda_{ij}=0$ if $i>j$.
@@
We start the proof by discussing the necessity of the above condition, to later switch to its sufficiency. Let us divide the populations of the Gibbs ensemble into two sets. The first are those corresponding to energies smaller of equal $E_{m}$ and the other correspond to larger energies. Since the operation $\Gamma$ is thermal, Gibbs ensemble is an invariant state, $\Gamma{\gamma}={\gamma}$, so the flow of populations between those two sets, described by $\Gamma$ elements must balance, as shown below
$\sum_{i>m}\sum_{j\leq m}\Lambda_{ij}{\gamma}_{j}=\sum_{i>m}(\sum_{j}\Lambda_{ij}{\gamma}_{j}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})=$ (43)
$=\sum_{i}({\gamma}_{i}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})-\sum_{i\leq m}({\gamma}_{i}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})=1-\sum_{j>m}{\gamma}_{j}-\sum_{i\leq m}{\gamma}_{i}+\sum_{i\leq m}\sum_{j<m}\Lambda_{ij}{\gamma}_{j}$
$=\sum_{i\leq m}\sum_{j<m}\Lambda_{ij}{\gamma}_{j}$
* To obtain the bounds on $\Lambda_{ij}$ elements, we may extrapolate the populations of Gibbs states using two inequalities: ${\gamma}_{i}\geq{\gamma}_{m}=e^{-\beta E_{m}}$ for $i\leq m$, and ${\gamma}_{i}\leq{\gamma}_{m+1}=e^{-\beta E_{m+1}}$ for $i>m$. Thus the above equality can be changed into
@@
$\sum_{i>m}\sum_{j\leq m}\Lambda_{ij}$ $\leq(\sum_{i\leq m}\sum_{j>m}\Lambda_{ij})e^{-\beta(E_{m+1}-E_{m})}.$
@@
It turns out, that, the condition from this Lemma is also sufficient. This can be seen by considering the channel, which corresponds to a detailed balance of Gibbs state, ie $\Lambda_{ij}{\gamma}_{j}=\Lambda_{ji}{\gamma}_{i}$, without any summation involved. Each such stochastic channel is by its construction also thermal. However, in the limit of zero temperature $\beta=\infty$, the condition of detained balance $\Lambda_{ij}=\Lambda_{ji}e^{-\beta(E_{i}-E_{j})}$, is always satisfied given $\Lambda_{ij}=0$ if $i>j$. Hence the freedom of all nonvanishing $\Lambda$ elements is restricted only by stochasticity.

### v2

@@ 5.4 Majorization-like restrictions on outputs of (bi)thermal tensors
Figure 8: The allowed region for ourput state $\mathbf{r}=T(\mathbf{p},\mathbf{q})$ in local dimention $d=3$. The two input states are denoted by black dots. All images of $T(\mathbf{p},\mathbf{q})$ are convex combinations of extremal images, denoted by coloured dots, thus they lie inside the red border. The green region is the upper bound on states accessible via bithermal operation according to Theorem 13. (a) In infinite temperature $\beta=0$ images via permutation tensors and other extremal tensors are represented by red and blue dots, respectively. The grey dots correspond to the images of $\mathbf{p}$ and $\mathbf{q}$ under permutation pre-processing, which does not affect the allowed region. (b) In zero temperature, $\beta\rightarrow\infty$ images via extremal tensors are denoted in red.
6 Examples and properties of LTOCC protocols
In this section we will provide two examples of application of LTOCC and their extensions. In the first subsection we will consider the problem of approximating basic nonlocal logic gates – CNOT and SWAP – using LTOCC with memory restricted to each round of the protocol. Then, we will consider the case of parallel LTOCC and possibility of generating correlations between the subsystems.
6.1 Approximating logic gates using LTOCC
In this subsection, we provide examples of LTOCC maps without memory, which approximate standard logical operations on bits: CNOT and SWAP (between subsystems). Since both of those operations are nonlocal, they cannot be constructed from just local thermal operations. On the other hand, both of those operations have a straightforward construction using LOCC, which we present.
To compare the original gate $M$ with the thermal counterpart $M^{(\beta)}$ we treat them as a (rescaled) probability distribution and calculate a total variation distance between them $\delta(M,M^{(\beta)})=\frac{1}{2}\sum_{ijkl}|M_{ij,kl}-M_{ij,kl}^{(\beta)}|$ [66]. Let us denote Alice’s and Bob’s inverse temperature as $\beta^{(A)}$ and $\beta^{(B)}$. For convenience let us also define $g_{A}=e^{-\beta^{(A)}(E_{1}^{(A)}-E_{0}^{(A)})}$, $g_{B}=e^{-\beta^{(B)}(E_{1}^{(B)}-E_{0}^{(B)})}$ where $E_{0}^{(A)},E_{1}^{(A)}$, $E_{0}^{(B)},E_{1}^{(B)}$ are the energies of ground state and excited states of Alice’s and Bob’s systems respectively.
We start with the construction of thermal CNOT by the following scheme. First Alice measures her bit and sets the information to Bob who depending on the outcome performs a (thermal) swap $S^{(\beta_{B})}$ between its energy levels. The resulting thermal approximant $CNOT^{(\beta)}$ has the following form:
$CNOT^{(\beta)}_{A\rightarrow B}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1-g_{B}&1\\ 0&0&g_{B}&0\end{pmatrix}$
Thus in the limit of large temperature $g_{B}\to 1$ the thermal approximant $CNOT^{(\beta)}$ coincides with the standard controlled NOT. The statistical distance between them is given by
$\delta(CNOT,CNOT^{(\beta)})=g_{B}$
The construction of thermal swap between subsystem $SWAP^{(\beta)}$, similarly to standard $SWAP$, requires three rounds of alternate thermal CNOT’s. Following this, one obtains two alternative approximations
$SWAP_{A}^{(\beta)}=CNOT^{(\beta)}_{A\rightarrow B}CNOT^{(\beta)}_{B\rightarrow A}CNOT^{(\beta)}_{A\rightarrow B}=\begin{pmatrix}1&0&0&0\\ 0&1-g_{A}&g_{B}&0\\ 0&g_{A}&\left(1-g_{B}\right){}^{2}&1-g_{B}\\ 0&0&\left(1-g_{B}\right)g_{B}&g_{B}\\ \end{pmatrix}$
$SWAP_{B}^{(\beta)}=CNOT^{(\beta)}_{B\rightarrow A}CNOT^{(\beta)}_{A\rightarrow B}CNOT^{(\beta)}_{B\rightarrow A}=\begin{pmatrix}1&0&0&0\\ 0&\left(1-g_{A}\right){}^{2}&g_{B}&1-g_{A}\\ 0&g_{A}&1-g_{B}&0\\ 0&\left(1-g_{A}\right)g_{A}&0&g_{A}\\ \end{pmatrix}$
which, in the limit of large temperatures $g_{A},g_{B}\to 1$, both coincide with standard $SWAP$
$SWAP=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\\ \end{pmatrix}\leavevmode\nobreak\ ,$
and the statistical distance between them is given by
$\delta(SWAP,SWAP_{A}^{(\beta)})=3-g_{A}-2g_{B}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \delta(SWAP,SWAP_{B}^{(\beta)})=3-2g_{A}-g_{B}.$
Thus, for any finite temperature, there is an inherent error associated with approximating even basic conditional logic operations when one attempts to implement them with operations preserving the thermal equilibria of the local systems.
6.2 Generating correlations from product input states
Finally, we return to the general discussion of one-round and two-round LTOCC between two parties to demonstrate its ability to generate correlations.
@@ 6.2 Generating correlations from product input states
First and foremost we stress that neither LTOCC+M nor $\mathcal{S}$LTOCC are (semilocal) thermal operations, contrary to the protocols without memory. Both of these protocols can induce correlations even on Gibbs ensemble $\gamma=\gamma^{(A)}\otimes\gamma^{(B)}$, which can be most clearly seen in the case of extremal operations in zero or infinite temperatures. However, before discussing these extreme scenarios let us first consider LTOCC without memory in a wider range of temperatures.
Observation 6.
Let $T_{jkl}$ be a thermal tensor, $\sum_{l}T_{jkl}\gamma_{l}=\gamma_{j}$, having an identity [or permutation matrix in infinite temperature case] as at least one layer $T_{jkl_{0}}=\delta_{j,k}$ or $T_{jkl_{0}}=\delta_{j,\sigma(k)}$ for some permutation $\sigma$]. Then the pair of states $\mathbf{p},\mathbf{q}$, second of which is a sharp state $q_{l}=\delta_{l,l_{0}}$ can be mapped into the maximally correlated state in one-round LTOCC.
Proof.
Let us take $\Lambda_{ik}=\delta_{i,k}$ (or $\Lambda_{ik}=\delta_{i,\sigma(k)}$ in infinite temperature) and define $M_{ij,kl}=\Lambda_{ik}T_{jkl}$, then the output distribution $\pi_{ij}$ is given by
$\pi_{ij}=\sum_{kl}M_{ij,kl}p_{k}q_{l}=\sum_{kl}\Lambda_{ik}T_{jkl}p_{k}\delta_{l,l_{0}}=\sum_{k}\delta_{i,k}\delta_{j,k}p_{k}=\delta_{i,j}p_{i}\leavevmode\nobreak\ ,$
and analogously for the extended case in the infinite temperature. Since output distribution is diagonal, by measuring one output, one knows about the result of the second output as well. Formally, while calculating conditional entropy between outputs one obtains
$H(A|B)=H(B|A)=-\sum_{ij}\pi_{ij}\ln\left(\frac{\pi_{ij}}{p_{i}}\right)=-\sum_{ij}\delta_{i,j}p_{i}\ln\left(\frac{p_{i}}{p_{i}}\right)=-\sum_{i}p_{i}\ln(1)=0.$ (43)
Notice that the requirement for thermal tensor demanded above can be guaranteed not only in zero or infinite temperature but is satisfied, for example, by construction (37). Additional discussion and examples of correlating properties of LTOCC in finite temperature are presented in the C.
Suppose we allow ourselves to use one round of parallel LTOCC, especially $\mathcal{S}$LTOCC, which is a subset of two rounds of LTOCC with memory. In that case the above observation can be extended to encompass non-sharp states, and even Gibbs state.
Observation 7.
@@
Consider a $\mathcal{S}LTOCC$ operation consisting of two copies of the same extremal (bi)thermal tensor $M_{ij,kl}=T_{ikl}T_{jkl}$, which entries are either 00 or $1$. Then for any product input state $\mathbf{p}\otimes\mathbf{q}$ the output state $\mathbf{\pi}=M(\mathbf{p}\otimes\mathbf{q})$ is maximally correlated.
@@
$\pi_{ij}=\sum_{kl}M_{ij,lk}p_{k}q_{l}=\sum_{kl}T_{ikl}T_{jkl}p_{k}q_{l}=\sum_{kl}\delta_{i,j}T_{ijk}p_{k}q_{l}=\delta_{ij}r_{i}\leavevmode\nobreak\ .$ (44)
Thus the output distribution is once again diagonal. ∎
The above requirements for thermal tensors are satisfied by extremal bithermal tensors in the limit of zero temperature $\beta\to\infty$ and a subset of extremal tensors in infinite temperature $\beta\to 0$ known as permutation tensors [56]. The above Observation points to the fact that LTOCC with long-term memory is a significantly more powerful framework than LTOCC without memory.
@@ 7 Summary and Outlook
In this work, we introduce a novel framework where parties operate under the distant laboratories paradigm, subject to thermodynamic constraints, coining the notion of local thermal operations and classical communication (LTOCC). We begin by presenting a hierarchy of classes of allowed operations, categorized by the level of control and the complexity of communication between the parties, and present the inclusion relations between them, including a special case of parallel LTOCC and its symmetric variant $\mathcal{S}$LTOCC. In particular, we show that LTOCC with shared randomness form a subset of semilocal thermal operations (SLTO), pointing to a possibility of realising finite-size thermal engines operating at Carnot efficiency, as described in [33], within the distant laboratory setting subject to classical communication restrictions. To do so, we provide new results for SLTO, presenting, among others, that the set of such operations is convex and closed under composition.
Since thermal tensors, encoding conditional thermal operations when restricted to the energy-incoherent states, constitute basic building blocks of the LTOCC framework, the second part of the work is devoted to the study of their properties. In particular, a major part of our considerations is concerned with the bithermal tensors, which constitute a symmetric subset of thermal tensors. The structure of the full set of such tensors can be understood as a thermal extension of the Birkhoff polytope of multistochastic tensors, and as such is an interesting mathematical problem in and of itself.
Finally, we consider application of LTOCC protocols to specific scenarios, starting with the task of approximating standard logical operations, CNOT and SWAP, while maintaining local equilibria of the systems under consideration. This serves to demonstrate the limitations of such operations – even when restricted to the energy-incoherent regime, LTOCC impose limits on how accurately one can perform logical gates. The second example considers possibility of generating correlations using LTOCC and their symmetric variant, demonstrating that under certain conditions it is possible to generate perfect correlations using either of the protocols.
@@
The results presented here lay the foundation for a new framework that combines two established concepts: the distant laboratories setting and thermodynamic limitations on operations. Given the novelty of this framework, several key questions remain open for future research. The first is a deeper exploration of multi-round scenarios, which may allow for a broader range of accessible state transitions, and formulation of stronger relations between protocols with different numbers of rounds. Another important question is the introduction of coherences into the framework. This is a challenging problem, as coherent thermal operations are difficult to describe in closed form even for a single qubit [67, 43]. The distinction between thermal operations and Gibbs-preserving operations, which emerge on the quantum level, further complicates this analysis. However, taking into account the conjecture that LTOCC with shared randomness and SLTO are equal when restricted to energy-incoherent processes, extension to coherent states in future work may provide tools necessary to identify the gap between classical and genuinely quantum effects in thermodynamics of finite-size systems interacting with multiple thermal baths.
@@
Moreover, in the face of recent developments at the frontier of quantum thermodynamics and nonlocality researchers within the community have begun to investigate interplays between characteristic features of both subfields, such as possibility of generating entanglement using out-of-equilibrium states [68], detection of nonlocal properties of states using their thermodynamic properties [69] or ever-growing body of literature concerned with steady-state entanglement from autonomous machines operating in thermal framework [70, 71]. In all of the above cases, we believe that considering the problem within the LTOCC framework would further understanding of both thermodynamical and informational tradeoffs incurred throughout the implementation of the protocol.
@@
All of these questions, while significant, fall outside the scope of this initial presentation of the LTOCC framework. They should be viewed as a roadmap for future studies, aimed at expanding and refining this new theoretical approach, and some of them will be explored in a future manuscript [61].
@@ Acknowledgments
RB acknowledges support by the National Science Centre, Poland, under the contract number 2021/03/Y/ST2/00193 within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733. JCz is supported by the start-up grant of the Nanyang Assistant Professorship at the Nanyang Technological University in Singapore, awarded to Nelly Ng.
@@ Appendix A Cooling maps
In this Appendix, we discuss the theoretical framework designed to grasp thermal operations in low temperatures: cooling maps [45]. Although the original construction was concerned with the general scenario of mixed quantum states, we let ourselves limit the discussion to diagonal states, similar to the bulk of the paper.
@@
Therefore the energy-preserving interaction with such a bath (3) can only deexcite the system of interest. Indeed, if we consider a diagonal input state with populations $\mathbf{p}=\operatorname{diag}_{H}(\rho)$, the action of the cooling map does not create coherences, as in the case of thermal operations, and modify the populations according to [45]
@@ Appendix B Alternative derivation of cooling maps
In this Appendix, we present an alternative derivation of cooling maps in the case of energy-incoherent states as a thermal operation in the limit of zero temperature $\beta\to\infty$.
@@
In the zero temperature limit $\beta\to\infty$, assuming nondegenerated Hamiltonian spectrum, the stochastic matrix $\Lambda$ corresponds to thermal operation if and only if $\Lambda_{i,j}=0$ for $i>j$. Furthermore, the extremal operations correspond to matrices with 00 and $1$ entries, restricted only by the condition $\Lambda_{ij}=0$ if $i>j$.
@@
We start the proof by discussing the necessity of the above condition, to later focus on its sufficiency. Let us divide the populations of the Gibbs ensemble into two sets. The first are those corresponding to energies smaller of equal $E_{m}$ and the other correspond to larger energies. Since the operation $\Lambda$ is thermal, Gibbs ensemble is an invariant state, $\Lambda{\gamma}={\gamma}$, so the flow of populations between those two sets, described by $\Gamma$ elements must balance, which can be shown explicitly,
$\sum_{i>m}\sum_{j\leq m}\Lambda_{ij}{\gamma}_{j}=\sum_{i>m}\quantity(\sum_{j}\Lambda_{ij}{\gamma}_{j}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})=$ (48)
$=\sum_{i}\quantity({\gamma}_{i}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})-\sum_{i\leq m}\quantity({\gamma}_{i}-\sum_{j>m}\Lambda_{ij}{\gamma}_{j})=1-\sum_{j>m}{\gamma}_{j}-\sum_{i\leq m}{\gamma}_{i}+\sum_{i\leq m}\sum_{j<m}\Lambda_{ij}{\gamma}_{j}$
$=\sum_{i\leq m}\sum_{j<m}\Lambda_{ij}{\gamma}_{j}\leavevmode\nobreak\ .$
To obtain the bounds on $\Lambda_{ij}$ elements, we may extrapolate the populations of Gibbs states using two inequalities: ${\gamma}_{i}\geq{\gamma}_{m}=e^{-\beta E_{m}}$ for $i\leq m$, and ${\gamma}_{i}\leq{\gamma}_{m+1}=e^{-\beta E_{m+1}}$ for $i>m$. Thus the above equality can be changed into
@@
$\sum_{i>m}\sum_{j\leq m}\Lambda_{ij}$ $\leq\quantity(\sum_{i\leq m}\sum_{j>m}\Lambda_{ij})e^{-\beta(E_{m+1}-E_{m})}.$
@@
It turns out, that, the condition from this Lemma is also sufficient. This can be seen by considering the channel, which corresponds to a detailed balance of Gibbs state, ie $\Lambda_{ij}{\gamma}_{j}=\Lambda_{ji}{\gamma}_{i}$, without any summation involved. Each such stochastic channel is by its construction also thermal. However, in the limit of zero temperature $\beta=\infty$, the condition of detailed balance $\Lambda_{ij}=\Lambda_{ji}e^{-\beta(E_{i}-E_{j})}$, is always satisfied given $\Lambda_{ij}=0$ if $i>j$. Hence the freedom of all nonvanishing $\Lambda$ elements is restricted only by stochasticity.